A certain airplane has two independent alternators to provide electrical power. The probability that a given alternator will fail on a 1-hour is 0.2. What is probability that (a) both will fail? (b) Neither will fail? (c ) One or the other will fail? Show all steps carefully.
Answer by stanbon(75887)

A certain airplane has two independent alternators to provide electrical power. The probability that a given alternator will fail on a 1-hour is 0.2. What is probability that
(a) P(both will fail) = 0.2^2 = 0.04
(b) P(Neither will fail) = (0.8)^2 = 0.16
(c) P(One or the other will fail) = 1 - [0.04+0.64] = 1 - 0.68 = 0.32
======================================
Cheers,
Stan H.